OpenStudy (anonymous):

Please help. I've been working on this for hours/: Let f(x)=x^2+6 and g(x)=(x+8)/x. Find (g o f)(-7) a) -55/7 b) 384/7 c) 295/49 d) 63/55 ANY help would be VErY much appreciated .(:

OpenStudy (anonymous):

find f(7) first then what you get for that you put it in as x of g(x)

OpenStudy (anonymous):

\[f(-7)=(-7)^2+6=55\] now find \(g(55)\)

OpenStudy (anonymous):

\[g(55)=\frac{55+8}{55}\] etc

OpenStudy (anonymous):

Wait what? I'm confused.....

OpenStudy (anonymous):

Plut -7 into the f(x) function first.

OpenStudy (anonymous):

Ok. so f(-7)=(-7)^2+6=55 rigt?

OpenStudy (anonymous):

yes, not you take the 55 and put it into the g(x) function

OpenStudy (anonymous):

So it's g(55)=(55+8)/x

OpenStudy (anonymous):

good, but also put 55 in for the x on the bottom.

OpenStudy (anonymous):

Oh right. My bad. So g(55)=(55+8)/55 = g(55)=63/55 ?

OpenStudy (anonymous):

there you go

OpenStudy (anonymous):

So the answer is 63/55?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thank you so much!!

OpenStudy (anonymous):

Would you be able to help me with one more?

OpenStudy (anonymous):

i can try